solve for the indicated variable
1/a + 1/b = c solve for a

Answer by user_dude2008(1862)

1/a + 1/b = c

b/ab + a/ab = c

(b+a)/ab = c

b+a= abc

b= abc - a

b=a(bc-1)

b/(bc-1)=a

Ans: a=b/(bc-1)